# RealMAN: A Real-Recorded and Annotated Microphone Array Dataset for Dynamic Speech Enhancement and Localization

### v1

@@ 4.3 Variable-array networks and array generalization
End-to-end speech enhancement and source localization networks are normally array-dependent, which means although the previous networks trained using real speech and real noise achieve better performance for one array, they still cannot be used for other arrays. In this section, we use all the 28 microphones data (microphone 0 $\sim$ 27) on the horizontal plane to train the variable-array networks, i.e. FaSNet-TAC [10] for speech enhancement and IPDnet [12] for sound source localization, to see whether the trained networks can be used as universal networks that can be directly used to unseen arrays.
@@
unprocessed 1.15 -9.8 2.00 1.72 1.51 19.9 1.11 -9.1 1.79 1.54 1.36 23.8
Fixed-Array 1.46 0.9 2.76 3.32 2.32 25.0 1.39 0.6 2.68 3.28 2.24 30.4
Variable-Array 1.40 -0.0 2.70 3.20 2.22 27.4 1.35 0.2 2.63 3.17 2.15 33.4
We set one test array, namely a 5-channel uniformly-spaced linear array (microphone 11, 3, 0, 7, and 12 in Fig. 1(b)). The training of variable-array networks uses randomly selected 2 $\sim$ 8-channel sub-arrays, excluding all 5-channel uniformly-spaced linear arrays.
Table 7 and Table 6 presents the results of speech enhancement and sound source localization, respectively. It can be seen that, there are indeed certain performance losses when compared with the fixed-array networks that are trained using the test array, but the losses are relatively small. This show that the 32-channel real-recorded microphone array data provided in the proposed dataset can successfully train the variable-array networks, which offers a competitive solution to the simu-to-real problem of multi-channel speech enhancement and sound source localization.
@@ 5 Conclusion
This paper presents a new real-recorded and annotated microphone array speech and noise dataset, called RealMAN, for speech enhancement and localization. Baseline experiments demonstrate that training with our real-recorded data outperforms training with simulated data, effectively eliminating the simulation-to-real gap for speech enhancement and localization. The performance on our dataset can better reflect the capabilities of tested algorithms in real-world applications, providing a more reliable benchmark for speech enhancement and localization. Additionally, unified variable-array networks can be successfully trained using various sub-arrays of the proposed 32-channel microphone array. These trained variable-array networks have the potential to be applied directly to unseen arrays in real-world applications.
@@
      Did you describe the limitations of your work? [Yes] See Sec. 2. One unreal factor is that we use a loudspeaker playing back speech signals, instead of speaking by real human speakers.
@@
      Did you discuss any potential negative societal impacts of your work? [Yes] There are no potential negative societal impacts. See supplementary materials.
@@ Appendix A Existing Microphone Array Speech Datasets
LibriCSS [20] 1 Meeting Static 10 h Recording, Transcription 7-ch circular
@@
CHiME-5/-6/-7 [22] 20 Home Moving 50 h Recording, Transcription Kinect ($\times$6), binaural pairs ($\times$4)
AMIMeeting [23] 3 Meeting - 100 h Recording, Transcription 8-ch circular, 8/4/10-ch
AISHELL-4 [24] 10 Meeting Static+ slightly moving 120 h Recording, Transcription 8-ch circular
AliMeeting [25] 21 Meeting Static 120 h Recording, Transcription 8-ch circular
@@ B.1 Scene information
We present detailed information of acoustic scenes for RealMAN recording in Table 9. The scene names in this table are consistent with the names in our released datasets. The reverberation time, i.e. T60, of indoor and semi-outdoor scenes are provided. The T60s of most enclosed scenes are measured using our recording system with the exponential sine sweep signal (see details in Appendix B.2). For some scenes that we cannot move our devices into, the T60s are measured by a mobile phone application, which however could be inaccurate.
@@
Table 9: Summary of recording scene information. The T60s of indoor scenes are measured by ourselves (see Appendix B.2 for measurement details), except that the underlined ones are measured by a less precise phone application.
Scene name Dataset Speech Noise T60 Category Room Size (m)
BarberShop Train $\times$ 0.653 Indoor -
Bus-Electric Train 0.283 Transportation -
Cafeteria2 Train $\times$ 1.242 Indoor 13.3$\times$9.2$\times$3.7
ConstructionPlant Train $\times$ - Outdoor -
LivingRoom3 Train $\times$ 0.376 Indoor 8.5$\times$5.3$\times$2.6
LivingRoom7 Train $\times$ 0.398 Indoor 12.9$\times$8.2$\times$2.2
LivingRoom9 Train $\times$ 0.562 Indoor 9.7$\times$9.6$\times$3.0
Park Train $\times$ - Outdoor -
PedestrianStreet Train $\times$ 0.336 Semi-outdoor -
Roadside1 Train $\times$ - Outdoor -
Roadside2 Train $\times$ - Outdoor -
ShoppingMall Train $\times$ 0.946 Indoor -
Subway-Electric Train $\times$ - Transportation -
Terrace1 Train $\times$ 0.355 Semi-outdoor -
Terrace2 Train $\times$ 0.605 Semi-outdoor -
UndergroundParking1 Train $\times$ 2.923 Indoor -
Auditorium Train,Validation $\times$ 1.179 Indoor -
BadmintonCourt1 Train,Validation 1.577 Indoor 28.6$\times$7.7$\times$8.6
Gym Train,Validation 1.504 Indoor 39.5$\times$8.5$\times$4.4
LivingRoom6 Train,Validation $\times$ 0.398 Indoor 3.6$\times$2.0$\times$2.6
LivingRoom8 Train,Validation $\times$ 0.399 Indoor 3.8$\times$2.0$\times$2.6
Market Train,Validation 1.228 Indoor 31.0$\times$30.0$\times$4.9
Cafeteria1 Train,Test 0.763 Indoor -
Car-Gasoline Train,Test 0.069 Transportation -
Classroom2 Train,Test $\times$ 0.679 Indoor 20.0$\times$16.4$\times$3.5
Classroom3 Train,Test $\times$ 1.358 Indoor 23.8$\times$6.9$\times$3.9
Library Train,Test $\times$ 1.180 Indoor radius=8.8
LivingRoom1 Train,Test 0.598 Indoor 9.7$\times$9.6$\times$3.0
LivingRoom2 Train,Test $\times$ 0.844 Indoor 4.5$\times$2.8$\times$2.9
LivingRoom4 Train,Test $\times$ 0.427 Indoor 7.7$\times$3.0$\times$2.0
LivingRoom5 Train,Test $\times$ 0.562 Indoor 4.2$\times$3.6$\times$2.9
OfficeRoom4 Train,Test $\times$ 0.888 Indoor 6.8$\times$5.1$\times$3.3
BasketballCourt1 Train,Validation,Test - Outdoor -
Car-Electric Train,Validation,Test 0.086 Transportation -
Classroom1 Train,Validation,Test* 0.816 Indoor 12.7$\times$7.6$\times$3.5
Laundry Train,Validation,Test $\times$ 0.534 Indoor 3.6$\times$2.6$\times$2.5
OfficeRoom3 Train,Validation,Test 1.295 Indoor 8.5$\times$6.0$\times$3.5
OfficeLobby Train,Validation,Test 2.476 Indoor -
SunkenPlaza1 Train,Validation,Test $\times$ 0.915 Semi-outdoor 27.6$\times$14.6
SubkenPlaza2 Train,Validation,Test $\times$ 0.915 Semi-outdoor -
BasketballCourt2 Validation - Outdoor -
Cafeteria3 Validation 0.843 Indoor -
OfficeRoom1 Validation $\times$ 0.719 Indoor 18.6$\times$10.5$\times$5.0
BadmintonCourt2 Test 1.694 Indoor 28.6$\times$7.7$\times$8.6
OfficeRoom2 Test 0.490 Indoor -
UndergroundParking2 Test 4.928 Indoor -
* The speech data of Classroom1 is split to the training set; And the corresponding noise data is split to train, validation and test sets.
B.2 T60 measurement
@@
The T60 is measured based on the RIR measurement with an exponential sine sweep signal. Firstly, we generate repeated exponential sine sweeps with a frequency ranging from 200 Hz to 8 kHz and their inverse signals to measure the RIR [41]. Secondly, the energy decay curve of the measured RIR is calculated. The appropriate segment in the energy decay curve is selected to calculate the T20 using linear least-squares regression. The T60 is approximated as three times the T20. For one scene, we conduct multiple trials to measure the T60, and the mean value is taken as the final T60 measurement.
@@ C.1 Details for target direct-path clean speech estimation
It is difficult to accurately estimate $A$ and $\tau$ from noisy and reverberant speech recordings, due to the signal distortion caused by reverberation and noise. Therefore, we leverage a prior speech enhancement step (using the AlignSpatialNet presented in Appendix C.2) which gives a rough estimation of the direct-path speech, denoted as $\hat{s}_{dp}(t)$, which is supposed to be roughly time- and level-aligned with ${s}_{dp}(t)$. Note that $\hat{s}_{dp}(t)$ is not used as the wanted target clean speech directly, because its interpretability and accuracy are not good enough. Alternatively, we calculate the generalized cross-correlation (GCC) [42] between $h_{dev}*s(t)$ and $\hat{s}_{dp}(t)$ to estimate the time shift $\tau$ (at the sample level) by finding the maximum peak of GCC. For static sources, $A$ and $\tau$ are time-invariant, and we estimate them at the utterance level. As for moving sources, $A$ and $\tau$ are time-varying. Because GCC is a statistical value, the estimation at the sample level is intractable. Therefore, we clip one speech utterance into overlapping segments and estimate $\tau$ at the segment level. The segment length of direct-path speech is determined by the time delay estimate of consecutive segments, and then the segments of source speech are stretched or compressed (by resampling operation) to align with the segments of direct-path speech. After that, the estimation of segment-level $A$ can also be done. Note that, data cleansing and nonlinear interpolation operations are adopted in both $\tau$ and $A$ estimations. After all, $\tau$ and $A$ estimates for moving sources are at the segment and sample levels, respectively. The details for target clean speech estimation of static and moving sources are shown in Algorithm 1 and 2, respectively. For simplicity, the sampling rate of the speech signal is marked in the parentheses and the downsampling operation is omitted.
@@ C.2 Network architecture for source-signal guided direct-path speech estimation
We propose a network named AlignSpatialNet for source-signal guided direct-path speech estimation. As shown in Fig. 3, the network has two branches, one for the recorded microphone signals and one for the played source clean speech signal. Each branch is extended from our previously proposed SpatialNet [28]. Besides the neural components from SpatialNet, i.e. the cross-band block, the multi-head self-attention module (MHSA) and the time-convolutional feedforward network (T-ConvFFN), each branch also contains multi-head cross-attention module (MHCA) for extracting the spatial cues, i.e. time shift and level attenuation, between the played source signal and the recorded microphone signals.
@@
where $\textbf{h}_{q}$, $\textbf{h}_{k}$ and $\textbf{h}_{v}$ are respectively the query, key and value vectors, which correspond to $\textbf{h}_{s}$, $\textbf{h}_{m}$ and $\textbf{h}_{m}$ in source branch and correspond to $\textbf{h}_{m}$, $\textbf{h}_{s}$ and $\textbf{h}_{s}$ in recording branch. $f$ and $t$ denote the frequency index and frame index, and $l$ is a preset value for the maximum number of frame shifts in one layer, which is set to a value corresponding to 200 ms in our experiment.
@@
C.3 Details for LED light detection with eyefish camera

### v2

@@ 4.3 Variable-array networks and array generalization
End-to-end speech enhancement and source localization networks are normally array-dependent, which means although the fixed-array networks (presented in the previous section) trained using real speech and real noise achieve better performance for one array, they still cannot be used for other arrays. In this section, we use all the 28-microphone data (microphone 0 $\sim$ 27) on the horizontal plane to train the variable-array networks, i.e. FaSNet-TAC [13] for speech enhancement and IPDnet [15] for source localization, to see whether the trained networks can be directly used to unseen arrays.
@@
unprocessed 1.14 -9.9 2.01 1.73 1.52 20.1 1.10 -9.0 1.80 1.54 1.37 23.7
Fixed-Array 1.41 1.5 2.76 3.25 2.28 25.5 1.36 0.5 2.72 3.18 2.23 30.0
Variable-Array 1.42 1.0 2.70 3.33 2.28 27.4 1.36 0.3 2.64 3.27 2.20 31.4
We set one test array, namely a 5-channel uniformly-spaced linear array (microphone 11, 3, 0, 7, and 12). The training of variable-array networks uses randomly selected 2 $\sim$ 8-channel sub-arrays, excluding all 5-channel uniformly-spaced linear arrays. The microphone 0 is always used and taken as the reference channel.
Table 7 and Table 6 present the results of speech enhancement and sound source localization, respectively. It can be seen that there are indeed certain performance losses when compared with the fixed-array networks that are trained using the test array, but the losses are relatively small. This shows that the 32-channel real-recorded microphone array data provided in the proposed dataset can successfully train the variable-array networks, which offers a competitive solution for real-world multi-channel speech enhancement and sound source localization.
@@ 5 Conclusion
This paper presents a new real-recorded and annotated microphone array speech and noise dataset, named RealMAN, for speech enhancement and source localization. Baseline experiments demonstrate that training with our real-recorded data outperforms training with simulated data, by eliminating the simulation-to-real gap. The performance on our dataset can better reflect the capabilities of tested algorithms in real-world applications, providing a more reliable benchmark for speech enhancement and source localization. Additionally, variable-array networks can be successfully trained using various sub-arrays of the proposed 32-channel microphone array, and they have the potential to be applied directly to unseen arrays in real-world applications.
@@
      Did you describe the limitations of your work? [Yes] See Sec. 1. One unreal factor is that we use a loudspeaker playing back speech signals, instead of speaking by real human speakers. Another limitation is that only an intermediate amount of Mandarin speech are used as source speech.
@@
      Did you discuss any potential negative societal impacts of your work? [Yes] There are no potential negative societal impacts.
@@ Appendix A Existing Microphone Array Speech Datasets
LibriCSS [23] 1 Meeting Static 10 hours Recording, transcription 7-ch circular
@@
CHiME-5/-6/-7 [25] 20 Home Moving 50 hours Recording, transcription Kinect ($\times$6), binaural pairs ($\times$4)
AMIMeeting [26] 3 Meeting - 100 hours Recording, transcription 8-ch circular, 8/4/10-ch
AISHELL-4 [27] 10 Meeting Static+ slightly moving 120 hours Recording, transcription 8-ch circular
AliMeeting [28] 21 Meeting Static 120 hours Recording, transcription 8-ch circular
@@ B.1 Scene information
We present detailed information of acoustic scenes and recordings for RealMAN in Table 9. The scene names in this table are consistent with the names in our released datasets. For each scene, the duration of static/moving speech and noise, the sound pressure level of speech/background (during speech recording) and noise, the reverberation time (T60) of indoor and semi-outdoor scenes, the room size when available are provided. The T60s of most enclosed scenes are measured using our recording system with the exponential sine sweep signal (see details in Appendix B.3). For some scenes that are not feasible for our T60 measurement, the T60s are measured by a mobile phone application, which however could be inaccurate.
B.2 Speech duration statistics across speakers
The speech duration statistics of different speakers in the dataset are given in Fig. 2.
Figure 2: Speech duration statistics across speakers in the RealMAN dataset. In the naming of speaker ID, the one beginning with ’P’ denotes speaker in read speech, and the other denotes speaker in free talk. The speaker ID is color-coded to distinguish the train, validation, and test sets.
@@ B.2 Speech duration statistics across speakers
Table 9: Summary of recording scene information. The T60s of indoor scenes are measured by ourselves (see Appendix B.3 for measurement details), except that the underlined ones are measured by a less precise phone application. The scene types are abbreviated as follows: semi. for semi-outdoor and trsp. for transportation.
Scene Duration of speech and noise in train/validation/test sets (minute) Sound pressure level (dB/dBA) T60 (s) Room Size (m)
name type static moving noise speech background noise
BasketballCourt1 Outdoor 34 / 10 / 39 40 / 8 / 42 315 / 52 / 158 68 / 60 63 / 47 69 / 53 - -
Car(Electric) Trsp. 51 / 42 / 23 - 172 / 129 / 129 71 / 63 61 / 35 77 / 48 0.09 -
OfficeRoom3 Indoor 68 / 3 / 5 81/ 7 / 8 214 / 72 / 72 70 / 63 56 / 29 63 / 44 1.30 8.5$\times$6.0$\times$3.5
OfficeLobby Indoor 47 / 29 / 5 51 / 28 / 8 144 / 42 / 21 64 / 57 56 / 38 61 / 50 2.48 -
SunkenPlaza1 Semi. 81 / 11 / 9 68 / 15 / 8 - 67 / 60 64 / 45 - 0.92 27.6$\times$14.6
BadmintonCourt1 Indoor 77 / 13 / - 40 / 7 / - 317 / 80 / - 61 / 55 50 / 27 70 / 58 1.58 28.6$\times$7.7$\times$8.6
Gym Indoor 73 / 1 / - 71 / 6 / - 293 / 33 / - 64 / 59 45 / 28 61 / 53 1.50 39.5$\times$8.5$\times$4.4
Market Indoor 90 / 9 / - 128 / 8 / - 255 / 29 / - 67 / 60 60 / 39 71 / 67 1.23 31.0$\times$30.0$\times$4.9
Auditorium Indoor 123 / 14 / - 73 / 11 / - - 62 / 56 50 / 32 - 1.18 -
LivingRoom6 Indoor 54 / 15 / - 51 / 18 / - - 69 / 64 47 / 28 - 0.40 3.6$\times$2.0$\times$2.6
LivingRoom8 Indoor 68 / 22 / - 49 / 14 / - - 73 / 68 54 / 33 - 0.40 3.8$\times$2.0$\times$2.6
Cafeteria1 Indoor 76 / - / 40 45 / - / 22 483 / - / 120 64 / 57 55 / 31 66 / 58 0.76 -
Car(Gasoline) Trsp. 100 / - / 17 - 189 / - / 81 69 / 64 58 / 34 92 / 63 0.07 -
Classroom2 Indoor 101 / - / 14 81 / - / 7 - 66 / 57 62 / 33 - 0.68 20.0$\times$16.4$\times$3.5
Classroom3 Indoor 86 / - / 39 55 / - / 26 - 68 / 61 55 / 43 - 1.36 23.8$\times$6.9$\times$3.9
Library Indoor 100 / - / 26 69 / - / 21 - 66 / 59 57 / 44 - 1.18 radius=8.8
LivingRoom2 Indoor 8 / - / 9 9 / - / 7 - 70 / 66 44 / 29 - 0.84 4.5$\times$2.8$\times$2.9
LivingRoom4 Indoor 200 / - / 36 147 / - / 26 - 70 / 63 56 / 36 - 0.43 7.7$\times$3.0$\times$2.0
LivingRoom5 Indoor 90 / - / 10 84 / - / 10 - 71 / 63 52 / 33 - 0.56 4.2$\times$3.6$\times$2.9
OfficeRoom4 Indoor 80 / - / 14 60 / - / 11 - 69 / 63 63 / 41 - 0.89 6.8$\times$5.1$\times$3.3
Bus(Electric) Trsp. 109 / - / - 7 / - / - 250 / - / - 71 / 65 64 / 43 78 / 63 0.28 -
Classroom1 Indoor 76 / - / - 72 / - / - 338 / 25 / 120 69 / 60 64 / 31 65 / 40 0.82 12.7$\times$7.6$\times$3.5
LivingRoom1 Indoor 81 / - / - 69 / - / - - / - / 7 68 / 62 48 / 32 68 / 60 0.60 9.7$\times$9.6$\times$3.0
LivingRoom3 Indoor 88 / - / - 99 / - / - - 69 / 62 54 / 34 - 0.38 8.5$\times$5.3$\times$2.6
LivingRoom7 Indoor 132 / - / - 84 / - / - - 67 / 61 59 / 38 - 0.40 12.9$\times$8.2$\times$2.2
UndergroundParking1 Indoor 125 / - / - 103 / - / - - 68 / 60 56 / 39 - 2.92 -
BasketballCourt2 Outdoor - / 33 / - - / 28 / - - / 166 / - 69 / 63 62 / 47 - - -
Cafeteria3 Indoor - / 34 / - - / 25 / - - / 286 / - 67 / 61 59 / 40 69 / 59 0.84 -
OfficeRoom1 Indoor - / 39 / - - / 36 / - - 70 / 64 65 / 34 - 0.72 18.6$\times$10.5$\times$5.0
BadmintonCourt2 Indoor - / - / 43 - / - / 28 - / - / 355 62 / 57 52 / 29 67 / 62 1.69 28.6$\times$7.7$\times$8.6
OfficeRoom2 Indoor - / - / 38 - / - / 37 - / - / 118 66 / 60 59 / 41 66 / 46 0.49 -
UndergroundParking2 Indoor - / - / 43 - / - / 27 233 / - / 100 67 / 62 59 / 44 67 / 51 4.93 -
BarberShop Indoor - - 213/ - / - - - 69 / 64 0.65 -
Cafeteria2 Indoor - - 443 / - / - - - 72 / 63 1.24 13.3$\times$9.2$\times$3.7
ConstructionPlant Outdoor - - 346 / - / - - - 69 / 58 - -
Laundry Indoor - - 28 / 7 / 12 - - 64 / 55 0.53 3.6$\times$2.6$\times$2.5
LivingRoom9 Indoor - - 45 / - / - - - 64 / 59 0.56 9.7$\times$9.6$\times$3.0
Park Outdoor - - 338 / - / - - - 73 / 59 - -
PedestrianStreet Semi. - - 495 / - / - - 73 / 61 0.34 -
Roadside1 Outdoor - - 169 / - / - - - 75 / 62 - -
Roadside2 Outdoor - - 163 / - / - - - 76 / 63 - -
ShoppingMall Indoor - - 327 / - / - - - 77 / 72 0.95 -
Subway(Electric) Trsp. - - 177 / - / - - - 86 / 71 - -
SunkenPlaza2 Semi. - - 167 / 36 / 36 - - 66 / 48 0.92 -
Terrace1 Semi. - - 239 / - / - - - 73 / 55 0.36 -
Terrace2 Semi. - - 26 / - / - - - 72 / 59 0.61 -
B.3 T60 and device impulse response measurement
@@
T60s are measured based on the RIR measurement with an exponential sine sweep signal. Firstly, we generate repeated exponential sine sweeps with a frequency ranging from 200 Hz to 8 kHz and their inverse signals to measure the RIR [50]. Secondly, the energy decay curve of the measured RIR is calculated. The appropriate segment in the energy decay curve is selected to calculate the T20 using linear least-squares regression. The T60 is approximated as three times the T20. For each scene, we conduct multiple trials to measure the T60, and the mean value is taken as the final T60 measurement.
For measuring the device impulse response (i.e. $h_{dev}$), we also first measure the RIR. Then any one clearly distinct path (say the direct path) in the RIR encodes the device impulse response, which is extracted as the measurement of $h_{dev}$.
B.4 Source location statistics
Fig. 3 shows the histogram of source azimuth, elevation and distance. The azimuth angles are mainly distributed in the range of 0 $\sim$ 180°. Sometimes, we need to put the microphone array close to one of the side walls for having a large source-to-array distance or not disturbing other people. To maintain the recording consistence, we decided to put the speaker mainly in the 0 $\sim$ 180°space even when the array is not put close to one side wall. Due to the symmetry of the array, we think the 0 $\sim$ 180°space is fairly sufficient for representing the whole space in terms of the evaluation of speech enhancement and source localization. The elevation angles are mainly distributed in the range of -5 $\sim$ 15°considering that normally the height of device and speaker are not distinct too much. The source-to-array distances are mainly distributed in the range of 0.5 m $\sim$ 5 m, covering the main range of most speech applications.
Fig.4 shows real examples of four typical types of speaker moving trajectory.
Figure 3: Histogram of speaker azimuth, elevation and distance.
(a) Axial walking
(b) Tangential walking
(c) Random walking
(d) Pacing
Figure 4: Four typical types of speaker moving trajectory. For each trajectory, the speaker height is constant, thus the moving trajectory is visualized only in the horizontal plane. Colors from lighter to darker indicate the time evolving.
B.5 SNR distribution of validation and test sets
The SNR distribution of the speech-noise mixed validation and test sets are shown in Fig. 5, whose mean values are $0.0$ dB and $-0.8$ dB, respectively.
Figure 5: Histogram of SNR for validation and test sets.
@@ C.1 Details for target direct-path clean speech estimation
It is difficult to accurately estimate $A$ and $\tau$ from noisy and reverberant speech recordings. Therefore, we leverage a prior speech enhancement step (using the AlignSpatialNet presented in Appendix C.2) which gives a rough estimation of the direct-path speech, denoted as $\hat{s}_{dp}(t)$, which is supposed to be roughly time- and level-aligned with ${s}_{dp}(t)$. Note that $\hat{s}_{dp}(t)$ is not directly used as the wanted target clean speech, because its interpretability and accuracy are not good enough. Alternatively, we calculate the generalized cross-correlation (GCC) [51] between $h_{dev}*s(t)$ and $\hat{s}_{dp}(t)$ to estimate the time shift $\tau$ (at the sample level) by finding the maximum peak of GCC. For static sources, $A$ and $\tau$ are time-invariant, and we estimate them at the utterance level. As for moving sources, $A$ and $\tau$ are time-varying. Because GCC is a statistical value, the estimation at the sample level is intractable. Therefore, we clip one speech utterance into overlapping segments and estimate $\tau$ at the segment level. The segment length of direct-path speech is determined by the time delay estimate of consecutive segments, and then the segments of source speech are stretched or compressed (by the resampling operation) to align with the segments of direct-path speech. After that, the estimation of segment-level $A$ can also be done. Note that, data cleansing and nonlinear interpolation operations are adopted for both $\tau$ and $A$ estimations. After all, $\tau$ and $A$ estimates for moving sources are at the segment and sample levels, respectively. The details for target clean speech estimation of static and moving sources are shown in Algorithm 1 and 2, respectively. For simplicity, the sampling rate of the speech signal is marked in the parentheses and the downsampling operation is omitted.
@@ C.2 Network architecture for source-signal guided direct-path speech estimation
We propose a network named AlignSpatialNet for source-signal guided direct-path speech estimation. As shown in Fig. 7, the network has two branches, one for the recorded microphone signals and one for the played source clean speech signal. Each branch is extended from our previously proposed SpatialNet [34]. Besides the neural components from SpatialNet, i.e. the cross-band block, the multi-head self-attention module (MHSA) and the time-convolutional feedforward network (T-ConvFFN), each branch also contains multi-head cross-attention module (MHCA) for extracting the spatial cues, i.e. time shift and level attenuation, between the played source signal and recorded microphone signals.
@@
where $\textbf{h}_{q}$, $\textbf{h}_{k}$ and $\textbf{h}_{v}$ are respectively the query, key and value vectors, which correspond to $\textbf{h}_{s}$, $\textbf{h}_{m}$ and $\textbf{h}_{m}$ in the source branch and correspond to $\textbf{h}_{m}$, $\textbf{h}_{s}$ and $\textbf{h}_{s}$ in the recording branch. $f$ and $t$ denote the frequency index and frame index, and $l$ is a preset value for the maximum number of frame shifts in one layer, which is set to a value corresponding to 200 ms in our experiment.
@@
C.3 Details for LED light detection with the fisheye camera